OpenStudy (anonymous):

solve for a 5/a+c/2b=7/3a

OpenStudy (anonymous):

-8/a=c/2b =>a=-4b/c

OpenStudy (anonymous):

not sure this is correct

OpenStudy (anonymous):

lets go slow. subtract \[\frac{5}{a}\] from both sides

OpenStudy (amistre64):

not really sure what the equation is either ;)